I am a parent trying to help solve 7th grade question-- haha- here it is,
9+6/x*6=21
I believe the answer is 3, but what is the formula used?- thanks

Answer by Alan3354(21608)

9+6/x*6=21
----------
Is it ((9+6)/x)*6 = 21 ?
or 9 + (6/x)*6 = 21 ?
or something else?